Please help me on the following problems- -(y+8)-5=4(y+2)-6y

Answer by stanbon(75887)

-(y+8)-5=4(y+2)-6y
------
-y-8-5 = 4y+8-6y
------
= -y-13 = -2y+8
----
y = 21
=============
Cheers,
Stan H.